Consider a product that consists of an assembly of 250 independent components. All of the components must be non-defective for the product to function satisfactorily. Calculate the proportions of products produced under Two Sigma and Six Sigma quality performances, respectively, that will be defective.
Answer #1

for 2 sigma process:

probability that a component will be good if process is two sigma =0.9545

hence probability that a component will be defective =1-P(all components are good) =1-(0.9545)250

=0.99999121

for 3 sigma process:

probability that a component will be good if process is three sigma =0.9973

hence probability that a component will be defective =1-P(all components are good) =1-(0.9973)250

=0.491308173